(1 point) If the differential equation m dx +4 dt + 7x = 0 dt2 is overdamped, the range of values for m is? Your answer willI AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

The differential equation of a damped harmonic oscillation is given as:

\small m{d^2x\over dt^2}+b{dx\over dt}+kx=0

where b is called the damping constant, and the oscillation is overdamped, underdamped, or critically damped according to:

  • Underdamped

    \small b^2<4km


  • Critically damped

    \small b^2=4km


  • Overdamped

    \small b^2>4km

Now, we can use the above conditions to solve our problem. We are given the equation

\small m{d^2x\over dt^2}+4{dx\over dt}+7x=0

Thus, here b = 4, k = 7

Now, the motion is given to be overdamped, hence we must have

\small b^2>4km \Rightarrow 16 > 7m \Rightarrow m < {16\over 7}

Thus, the range of values of m for which the motion will be underdamped is

\small \left (-\infty, {16\over 7}\right )

write (-inf,16/7) in the answering box